# FedQUIT: On-Device Federated Unlearning via a Quasi-Competent Virtual Teacher

### v2

@@ Abstract
Federated Learning (FL) systems enable the collaborative training of machine learning models without requiring centralized collection of individual data. FL participants should have the ability to exercise their right to be forgotten, ensuring their past contributions can be removed from the learned model upon request. In this paper, we propose FedQUIT, a novel algorithm that uses knowledge distillation to scrub the contribution of the data to forget from an FL global model while preserving its generalization ability. FedQUIT directly works on client devices that request to leave the federation, and leverages a teacher-student framework. The FL global model acts as the teacher, and the local model works as the student. To induce forgetting, FedQUIT tailors the teacher’s output on local data (the data to forget) penalizing the prediction score of the true class. Unlike previous work, our method does not require hardly viable assumptions for cross-device settings, such as storing historical updates of participants or requiring access to proxy datasets. Experimental results on various datasets and model architectures demonstrate that (i) FedQUIT outperforms state-of-the-art competitors in forgetting data, (ii) has the exact computational requirements as a regular FedAvg round, and (iii) reduces the cumulative communication costs by up to 117.6$\times$ compared to retraining from scratch to restore the initial generalization performance after unlearning.
@@ 1 Introduction
In 2016, the European Union introduced the General Data Protection Regulation (GDPR), incorporating guidelines designed to empower individuals with full control over their personal data [7]. One of the cornerstones of GDPR is the principle of data minimization, which involves collecting and processing only the personal data that is necessary for the intended purpose and limiting the retention of such data to the minimum necessary time period. Such a requirement is not easily verifiable once the data are outsourced from the data owners’ machines or devices. Especially in the case of datacenter training of Machine Learning (ML) or Deep Learning (DL) models, the collected data may be easily reused to train ML/DL models with scopes different from what was initially declared.
@@
Federated Learning (FL) [24], an alternative to traditional centralized ML, allows the training of models while minimizing the personal information disclosed by the users. In a nutshell, FL learns a shared global model by periodically gathering and aggregating ephemeral model updates locally computed on private data, thus avoiding the transfer of unprocessed data.
@@
The right to be forgotten is another pivotal requirement described in the GDPR. Specifically, Article 17 of GDPR states that the data subject shall have the right to obtain from the controller the erasure of personal data where the data subject withdraws consent on which the processing is based. However, when data are used to train DL models, simply deleting the samples for which the permission has been revoked is not enough. DL models can memorize and then leak private information contained in the training corpus, as widely demonstrated during the last years, e.g., via membership inference attacks (MIAs) [28, 30, 29]. Under this premise, Machine Unlearning (MU) [8, 34] has gained traction as a set of mechanisms to selectively remove the contribution of specific training samples on a learned model without destroying the generalization ability of such model. The straightforward solution would be to retrain the model, excluding the data that must be forgotten. Still, this approach can be extremely costly and sometimes unfeasible, particularly when the training dataset is no longer available. While in centralized machine learning, this naive retraining strategy might be a viable, though inefficient, option, it becomes impractical in FL. This impracticality arises due to clients’ sporadic participation, especially in cross-device federated settings [2, 16]. Consequently, developing unlearning mechanisms specifically designed for federated environments is crucial. Additionally, unlike MU, Federated Unlearning (FU) must consider that the so-called retain data, i.e., the data that remains in the process, is not directly accessible by design [27].
@@
No Historical No Full No Stateful No Proxy No Comp. Work Information Participation Clients Data Intensive Liu et al. [22] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ FedEraser [20] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ FUKD [32] $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ PGA [12] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ FedQUIT (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
In this paper, we present an unlearning method tailored for FL settings. Unlike recent works (e.g., [4, 20, 32, 12]), our solution is lightweight and practical for any federated client wishing to exit the process by removing its contribution. It requires no historical update storage, stateful clients, full client participation, or proxy data.
@@
  • FedQUIT can quickly recover the gold-standard generalization ability and significantly reduce the communication cost compared to a state-of-the-art baseline. FedQUIT achieves up to 117.6$\times$ communication cost reduction compared to retraining from scratch.
@@ 2 Related Work
Table 1 provides a comparison with existing client unlearning methods, emphasizing key design limitations. An extended explanation of the column definitions can be found in Appendix 10. The method in [22] enables rapid retraining from scratch using historical information and Quasi-Newton methods but incurs significant computational overhead. Moreover, note that: (1) linking per-client update histories to specific clients requesting unlearning undermines FL’s privacy design; (2) to some extent it violates the ephemeral updates requirement of the FL framework [16]; and (3) at scale, considering a massive number of clients [6], several FL rounds, and multiple simultaneous learning tasks, the corresponding storage requirements might become impractical. FedEraser [20], a pioneering client unlearning approach, recalibrates historical updates but still requires the server to store all past client updates. FUKD [32] involves two server-side steps: first, it removes the unlearned client’s updates from the global model to create a sanitized model; then, it uses the previous global model as a teacher in a KD framework to improve the sanitized model’s performance, employing unlabeled proxy data. However, this approach assumes access to unlabeled data on the server, which may not always be feasible, and again requires storing historical updates. The approach in [12] resembles ours in terms of the protocol, where unlearning is performed locally by the client. Halimi et al. apply projected gradient ascent (PGA) to eliminate the influence of the data to be forgotten, constraining local weight updates to ensure that the model weights remain within an $l2$-norm ball around a reference model. However, PGA requires clients to store their latest model updates (stateful clients) and requires full client participation, as the unlearning process begins with a reference model derived by removing the last unlearning client’s model update from the current global model. In addition, PGA needs to tune, at least, a threshold that measures the distance between the unlearned model and the reference model to early stop the unlearning. Other works such as FAST [11] and the work in [1] primarily focus on backdoor removal, and, respectively, use historical information and suppose to have the unlearning client dataset divided in benign and trigger subsets. Unlike previous methods, our solution does not require stateful clients, full participation, or access to proxy data on the server. Furthermore, our method does not need to store the entire history of per-client model updates and link them to the client that requests unlearning.
@@
Related work in centralized MU. The most similar work to ours in non-federated settings is the mechanism presented in [5], where the authors use an incompetent teacher on forget data and a competent teacher, i.e., the model before unlearning, on retain data. However, this approach is not directly applicable in federated settings since retain data – clients’ data – are not accessible. For completeness, in the experimental results, we adapt our solution to centralized settings and compare it to [5], showing superior forget effectiveness and robustness to further fine tuning after forget.
@@ 3.1 Federated Learning
where $w$ represents the parameters of the global model to be trained, $n_{k}$ represents the size of $D_{k}$, and $n$ is the total number of examples held by all participating clients. $F_{k}$ is a local objective function. Federated Averaging (FedAvg) [24] is a synchronous client-server FL algorithm that operates in rounds of specific steps to solve the optimization problem of Eq. 1. At each round, the global model is distributed to a random subset of clients, which in turn locally fine-tune the model for $E$ epochs. Then, the activated clients send back a model update. The parameter server collects and aggregates the updates using weighted averaging based on the amount of per-client data samples. Finally, the server applies the averaged updates to the current global model [26].
@@ 3.2 Federated Unlearning
FU can have different objectives, i.e., sample unlearning, class unlearning [31], feature unlearning [9], and client unlearning. In this paper, we will consider the client unlearning case, where $S_{u}=D_{u}$. In practice, a client $u$ requests that the contribution of its entire local dataset is removed from the global model. Similarly to regular MU, the union of the remaining clients’ dataset $D_{r}$ is termed as retain dataset, i.e., $D_{r}=\bigcup_{\forall k\neq u}D_{k}$. The problem of FU is to design an effective and efficient forgetting mechanism to produce $w^{\bar{u}}_{t}$ from $w_{t}$. While the efficiency of an FU method is usually measured as the amount of FL rounds needed to reach the generalization performance of the retrained model (e.g., the test accuracy or loss), the ways to assess forgetting efficacy are more varied, but usually compares the unlearned model’s performances on the forgetting data with the retrained model’s ones [27].
@@ 4.1 Preliminaries
[…]
@@ 4.2 Local Unlearning
This entails that the unlearning method has direct access only to the target client’s data (the forget data), while the rest of the data in the federation (the retain data) would not be available, except via some form of proxy. Access to the knowledge derived from retain data during the unlearning phase is crucial to preserve the generalization performance of the global model while selectively removing the contribution of the forget data. A natural proxy for the data in the federation is represented by the current global model, and in our proposed method, we exploit it. As we discuss in Appendix 9, previous work has used the global model as a guide locally at clients, for example, to preserve knowledge in the presence of catastrophic forgetting and heterogeneous data. Our novel family of unlearning algorithms, FedQUIT, leverages a KD-based [14] teacher-student framework, where the global model works as teacher and the local model works as student. To induce forgetting, FedQUIT tailors the output of the teacher on forget data – the local data – penalizing the prediction score of the true class (with various possible methods explored below in this Section). In this sense, we refer to such a crafted version of the global model, as quasi-competent virtual teacher. It is worth noting that, in standard FL algorithms such as FedAvg [24], the global model is always sent to the selected clients as the starting point for the training round, so actively using it for inference on local data does not imply any further assumption over the regular FL procedure. An extended discussion on the inspirations for the design of FedQUIT can be found in Appendix 8.
@@
4:$w^{\bar{u}}\leftarrow w$triangleright Client $u$ initializes unlearned model
@@
6:     for each minibatch =$\mathcal{B}=\{(x_{i},y_{i})\}_{i=1}^{B}\in D_{u}$ do
@@
FedQUIT-Softmax. FedQUIT-Softmax uses the same KD-based framework of FedQUIT-Logits, with the global model at round $t$ that works as the teacher and the unlearned model that works as the student. Differently from FedQUIT-Logit, FedQUIT-Softmax modifies the output of the global model (the teacher) after the softmax has been applied to its logits. As before, we indicate the output probabilities of the global model at round $t$ with $g(t)$. FedQUIT-Softmax replaces the true-class output probability $g(y)$ with a fixed value of $v$ and uniformly distributes the remaining quantity $g(y)-v$ to all the other class probabilities $g(c)$.
@@ 5.1 Settings, datasets and baselines
We conduct experiments on federated versions of two datasets, CIFAR-10 and CIFAR-100 [18]. To simulate a federated environment, we partition each dataset among 10 clients, ensuring no repeated samples. We consider both IID and non-IID data distributions. In the non-IID setting, we introduce data heterogeneity using distribution-based label skew, controlled via Latent Dirichlet Allocation [15], with concentration parameters of $\alpha=0.3$ for CIFAR-10 and $\alpha=0.1$ for CIFAR-100. Fig. 6 in the Appendix illustrates the label distribution across clients.
We employ a standard ResNet-18 [13] and a visual transformer, namely the MiT-B0 [33]. Unless stated otherwise, before performing unlearning, we train ResNet-18 from scratch for 200 FedAvg rounds while we fine-tune MiT-B0 for 50 FedAvg rounds, starting from a pre-trained checkpoint, with one local training epoch per round ($E=1$). Additionally, we validate our results using a higher number of local epochs ($E=10$) for the CIFAR-100/ResNet-18 setting with non-IID data ($\alpha=0.1$). Table 2 outlines the various settings of our experiments. For each setting, we conduct 10 experiments, each with a different client designated as the unlearning client, and we report the mean and standard deviation of the results.
@@
Setting Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. MIA [29] MIA [36] CIFAR-100, IID, $E$=1 Original 59.9±0.0 79.8±0.7 78.2±0.6 71.4 ±0.6 Retrain 58.3±0.5 58.0±0.7 55.6±0.7 48.8 ±0.6 PGA [12] 12.1 ±4.2 16.7$\times$ 59.1±0.6 72.8 (14.8 ±1.0) 71.8 (16.1±1.8) 63.0 (14.1±1.6) Incompetent T. 15.6 ±3.8 12.8$\times$ 58.6 ±0.2 68.8 (10.8 ±2.0) 67.6 (12.0 ±1.8) 52.2 (13.4 ±1.8) FedQUIT 10.0±1.4 20.0$\times$ 58.7 ±0.4 59.4 (1.4±0.8) 60.7 (5.1±1.6) 55.7 (6.8±0.8) CIFAR-100, Non-IID, $E$=1 Original 53.8±0.0 62.9±6.9 75.4±7.0 61.0 ±7.8 Retrain 51.0±1.4 33.5±4.5 44.0±5.5 32.0 ±5.2 PGA [12] 7.0±5.1 28.6$\times$ 51.4 ±0.6 37.4 (4.3±4.1) 49.2 (6.4±4.9) 36.6 (5.3±5.3) Incompetent T. 1.7±1.25 117.6$\times$ 52.05±0.82 48.41 (14.94±6.05) 61.03 (17.04±6.50) 45.39 (13.38±6.02) FedQUIT 3.7±2.9 54.1$\times$ 52.0±0.5 34.6 (3.5±2.4) 50.7 (5.5±2.8) 33.9 (3.8±2.8) CIFAR-100, Non-IID, $E$=10 Original 50.3±0.0 60.8±6.6 73.4±6.6 61.3 ±7.5 Retrain 48.0±0.8 31.6±4.5 42.1±5.2 31.2 ±4.9 PGA [12] 4.4±3.8 45.5$\times$ 49.0±1.0 37.3 (6.2±8.1) 49.6 (7.7±9.6) 37.7 (6.6±8.2) Incompetent T. 1.0 ±0.0 200.0$\times$ 49.6 ±0.5 48.0 (16.6 ±6.4) 60.6 (18.4 ±7.7) 45.7 (14.6 ±6.7) FedQUIT 1.7±0.5 117.6$\times$ 48.3 ±0.4 27.9 (3.4±2.4) 45.9 (4.7±2.8) 29.9 (4.1±2.9) CIFAR-10, Non-IID, $E$=1 Original 83.72 ±0.00 88.59 ±3.85 84.39 ±5.35 81.37 ±6.28 Retrained 83.49 ±1.57 81.14 ±8.00 73.12 ±13.68 72.49 ±10.78 PGA [12] 11.5±5.6 17.4$\times$ 83.99 ±1.02 84.33 (3.19 ±4.86) 78.50 (5.38 ±7.70) 77.01 (4.51 ±6.04) Incompetent T. 1.1 ±0.3 181.8$\times$ 84.78 ±0.94 88.18 (7.04 ±5.08) 81.36 (8.24 ±5.96) 81.98 (9.49 ±6.34) FedQUIT 3.8 ±1.9 52.6$\times$ 83.74 ±1.17 81.71 (2.43 ±2.69) 72.12 (3.45 ±2.43) 71.06 (2.82 ±1.50)
@@
We compare our method with the Incompetent Teacher baseline (Section 2) and the state-of-the-art PGA baseline [12]. To the best of our knowledge, PGA is the only approach that avoids impractical assumptions (see Table 2 for an extended comparison), such as requiring access to the full history of client updates, which necessitates linking clients to their past updates and compromises the privacy design of federated learning. For fairness, we enforce full client participation, as required by PGA, even though our method naturally supports the standard partial participation setting. For FedQUIT, we always use one single epoch for unlearning ($E_{u}=1$) at the target client. We report the results for FedQUIT implemented as FedQUIT-Logits with $v=0$, as it performed best. In Sec. 6.1, we also discuss the results of FedQUIT-Softmax and other choices of $v$ value. See Appendix 8 for hyper-parameter tuning details.

### v3

@@ Abstract
[…]
@@ 1 Introduction
In 2016, the European Union introduced the General Data Protection Regulation (GDPR), incorporating guidelines designed to empower individuals with full control over their personal data [GDPR2016a]. One of the cornerstones of GDPR is the principle of data minimization, which involves collecting and processing only the personal data that is necessary for the intended purpose and limiting the retention of such data to the minimum necessary time period. Such a requirement is not easily verifiable once the data are outsourced from the data owners’ machines or devices. Especially in the case of datacenter training of Machine Learning (ML) or Deep Learning (DL) models, the collected data may be easily reused to train ML/DL models with scopes different from what was initially declared.
@@
Federated Learning (FL) [mcmahan2017communication], an alternative to traditional centralized ML, allows the training of models while minimizing the personal information disclosed by the users. In a nutshell, FL learns a shared global model by periodically gathering and aggregating ephemeral model updates locally computed on private data, thus avoiding the transfer of unprocessed data.
@@
The right to be forgotten is another pivotal requirement described in the GDPR. Specifically, Article 17 of GDPR states that the data subject shall have the right to obtain from the controller the erasure of personal data where the data subject withdraws consent on which the processing is based. However, when data are used to train DL models, simply deleting the samples for which the permission has been revoked is not enough. DL models can memorize and then leak private information contained in the training corpus, as widely demonstrated during the last years, e.g., via membership inference attacks (MIAs) [shokri2017membership, song2019privacy, song2021systematic]. Under this premise, Machine Unlearning (MU) [golatkar2020eternal, xu2023machine] has gained traction as a set of mechanisms to selectively remove the contribution of specific training samples on a learned model without destroying the generalization ability of such model. The straightforward solution would be to retrain the model, excluding the data that must be forgotten. Still, this approach can be extremely costly and sometimes unfeasible, particularly when the training dataset is no longer available. While in centralized machine learning, this naive retraining strategy might be a viable, though inefficient, option, it becomes impractical in FL. This impracticality arises due to clients’ sporadic participation, especially in cross-device federated settings [bellavista2021decentralised, kairouz2019advances]. Consequently, developing unlearning mechanisms specifically designed for federated environments is crucial. Additionally, unlike MU, Federated Unlearning (FU) must consider that the so-called retain data, i.e., the data that remains in the process, is not directly accessible by design [romandini2024federated].
@@
No Historical No Full No Stateful No Proxy No Comp. Work Information Participation Clients Data Intensive Liu et al. [liu2022right] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ FedEraser [federaser] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ FUKD [wu2022unlearningdistillation] $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ PGA [halimi2022federated] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ FedQUIT (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
In this paper, we present an unlearning method tailored for FL settings. Unlike recent works (e.g., [fedrecover, federaser, wu2022unlearningdistillation, halimi2022federated]), our solution is lightweight and practical for any federated client wishing to exit the process by removing its contribution. It requires no historical update storage, stateful clients, full client participation, or proxy data.
@@
[…]
@@ 2 Related Work
Table 1 provides a comparison with existing client unlearning methods, emphasizing key design limitations. An extended explanation of the column definitions can be found in Appendix 10. The method in [liu2022right] enables rapid retraining from scratch using historical information and Quasi-Newton methods but incurs significant computational overhead. Moreover, note that: (1) linking per-client update histories to specific clients requesting unlearning undermines FL’s privacy design; (2) to some extent it violates the ephemeral updates requirement of the FL framework [kairouz2019advances]; and (3) at scale, considering a massive number of clients [dean2012large], several FL rounds, and multiple simultaneous learning tasks, the corresponding storage requirements might become impractical. FedEraser [federaser], a pioneering client unlearning approach, recalibrates historical updates but still requires the server to store all past client updates. FUKD [wu2022unlearningdistillation] involves two server-side steps: first, it removes the unlearned client’s updates from the global model to create a sanitized model; then, it uses the previous global model as a teacher in a KD framework to improve the sanitized model’s performance, employing unlabeled proxy data. However, this approach assumes access to unlabeled data on the server, which may not always be feasible, and again requires storing historical updates. The approach in [halimi2022federated] resembles ours in terms of the protocol, where unlearning is performed locally by the client. halimi2022federated apply projected gradient ascent (PGA) to eliminate the influence of the data to be forgotten, constraining local weight updates to ensure that the model weights remain within an $l2$-norm ball around a reference model. However, PGA requires clients to store their latest model updates (stateful clients) and requires full client participation, as the unlearning process begins with a reference model derived by removing the last unlearning client’s model update from the current global model. In addition, PGA needs to tune, at least, a threshold that measures the distance between the unlearned model and the reference model to early stop the unlearning. Other works such as FAST [guo2023FAST] and the work in [alam2023get] primarily focus on backdoor removal, and, respectively, use historical information and suppose to have the unlearning client dataset divided in benign and trigger subsets. Unlike previous methods, our solution does not require stateful clients, full participation, or access to proxy data on the server. Furthermore, our method does not need to store the entire history of per-client model updates and link them to the client that requests unlearning.
@@
Related work in centralized MU. The most similar work to ours in non-federated settings is the mechanism presented in [chundawat2023can], where the authors use an incompetent teacher on forget data and a competent teacher, i.e., the model before unlearning, on retain data. However, this approach is not directly applicable in federated settings since retain data – clients’ data – are not accessible. For completeness, in the experimental results, we adapt our solution to centralized settings and compare it to [chundawat2023can], showing superior forget effectiveness and robustness to further fine tuning after forget.
@@ 3.1 Federated Learning
where $w$ represents the parameters of the global model to be trained, $n_{k}$ represents the size of $D_{k}$, and $n$ is the total number of examples held by all participating clients. $F_{k}$ is a local objective function. Federated Averaging (FedAvg) [mcmahan2017communication] is a synchronous client-server FL algorithm that operates in rounds of specific steps to solve the optimization problem of Eq. 1. At each round, the global model is distributed to a random subset of clients, which in turn locally fine-tune the model for $E$ epochs. Then, the activated clients send back a model update. The parameter server collects and aggregates the updates using weighted averaging based on the amount of per-client data samples. Finally, the server applies the averaged updates to the current global model [reddi2020adaptive].
@@ 3.2 Federated Unlearning
FU can have different objectives, i.e., sample unlearning, class unlearning [wang2023classdiscriminative], feature unlearning [gu2025ferrari], and client unlearning. In this paper, we will consider the client unlearning case, where $S_{u}=D_{u}$. In practice, a client $u$ requests that the contribution of its entire local dataset is removed from the global model. Similarly to regular MU, the union of the remaining clients’ dataset $D_{r}$ is termed as retain dataset, i.e., $D_{r}=\bigcup_{\forall k\neq u}D_{k}$. The problem of FU is to design an effective and efficient forgetting mechanism to produce $w^{\bar{u}}_{t}$ from $w_{t}$. While the efficiency of an FU method is usually measured as the amount of FL rounds needed to reach the generalization performance of the retrained model (e.g., the test accuracy or loss), the ways to assess forgetting efficacy are more varied, but usually compares the unlearned model’s performances on the forgetting data with the retrained model’s ones [romandini2024federated].
@@ 4.1 Preliminaries
Considering a supervised classification task, for each $d$-dimensional input pattern $x\in\mathbb{R}^{d}$, a neural network $h(\cdot;w)$ with parameters $w$ produces the probabilities, denoted as $p(c|x)$, for each label $c\in\{1,...,C\}$ with $C$ representing the task classes. The output probabilities are obtained using the Softmax activation on the logits $z_{c}$ generated by $h(x;w)$, i.e., $p(c|x)=\frac{exp(z_{c}/\tau)}{\sum_{j=1}^{C}exp(z_{j}/\tau)}$ with $\tau$ being the temperature. The true distribution of labels for a given input $x$ is represented as $q(c|x)$. When using hard targets, with $y$ being the ground-truth label for sample $x$, we have $q(y|x)=1$, indicating that $y$ is the correct label, and $q(c|x)=0$ for all $c\neq y$, indicating that all the other labels are incorrect. To simplify notation, we use $p(c)$ to refer to $p(c|x)$ and $q(c)$ to refer to $q(c|x)$. The model $h(\cdot;w)$ is usually trained by minimizing the cross-entropy loss between the true targets $q(c)$ and the network’s outputs $p(c)$ as in $H(q,p)=\sum_{k=1}^{C}-q(c)log(p(c))$.
@@ 4.2 Local Unlearning
This entails that the unlearning method has direct access only to the target client’s data (the forget data), while the rest of the data in the federation (the retain data) would not be available, except via some form of proxy. Access to the knowledge derived from retain data during the unlearning phase is crucial to preserve the generalization performance of the global model while selectively removing the contribution of the forget data. A natural proxy for the data in the federation is represented by the current global model, and in our proposed method, we exploit it. As we discuss in Appendix 9, previous work has used the global model as a guide locally at clients, for example, to preserve knowledge in the presence of catastrophic forgetting and heterogeneous data. Our novel family of unlearning algorithms, FedQUIT, leverages a KD-based [hinton2015distilling] teacher-student framework, where the global model works as teacher and the local model works as student. To induce forgetting, FedQUIT tailors the output of the teacher on forget data – the local data – penalizing the prediction score of the true class (with various possible methods explored below in this Section). In this sense, we refer to such a crafted version of the global model, as quasi-competent virtual teacher. It is worth noting that, in standard FL algorithms such as FedAvg [mcmahan2017communication], the global model is always sent to the selected clients as the starting point for the training round, so actively using it for inference on local data does not imply any further assumption over the regular FL procedure. An extended discussion on the inspirations for the design of FedQUIT can be found in Appendix 8.
@@
[…]
@@
6:  for each minibatch $\mathcal{B}=\{(x_{i},y_{i})\}_{i=1}^{B}\in D_{u}$ do
@@
FedQUIT-Softmax. FedQUIT-Softmax uses the same KD-based framework of FedQUIT-Logits, with the global model at round $t$ that works as the teacher and the unlearned model that works as the student. Differently from FedQUIT-Logit, FedQUIT-Softmax modifies the output of the global model (the teacher) after the softmax has been applied to its logits. As before, we indicate the output probabilities of the global model at round $t$ with $g(t)$. FedQUIT-Softmax replaces the true-class output probability $g(y)$ with a fixed value of $v$ and uniformly distributes the remaining quantity g(y)vg(y)-v to all the other class probabilities $g(c)$.
@@ 5.1 Settings, datasets and baselines
We conduct experiments on federated versions of two datasets, CIFAR-10 and CIFAR-100 [Krizhevsky09learningmultiple]. To simulate a federated environment, we partition each dataset among 10 clients, ensuring no repeated samples. We consider both IID and non-IID data distributions. In the non-IID setting, we introduce data heterogeneity using distribution-based label skew, controlled via Latent Dirichlet Allocation [hsu2019measuring], with concentration parameters of $\alpha=0.3$ for CIFAR-10 and $\alpha=0.1$ for CIFAR-100. Fig. 6 in the Appendix illustrates the label distribution across clients.
We employ a standard ResNet-18 [he2016deep] and a visual transformer, namely the MiT-B0 [xie2021segformer]. Unless stated otherwise, before performing unlearning, we train ResNet-18 from scratch for 200 FedAvg rounds while we fine-tune MiT-B0 for 50 FedAvg rounds, starting from a pre-trained checkpoint, with one local training epoch per round ($E=1$). Additionally, we validate our results using a higher number of local epochs ($E=10$) for the CIFAR-100/ResNet-18 setting with non-IID data ($\alpha=0.1$). Table 2 outlines the various settings of our experiments. For each setting, we conduct 10 experiments, each with a different client designated as the unlearning client, and we report the mean and standard deviation of the results.
@@
[…]
@@
We compare our method with the Incompetent Teacher baseline (Section 2) and the state-of-the-art PGA baseline [halimi2022federated]. To the best of our knowledge, PGA is the only approach that avoids impractical assumptions (see Table 2 for an extended comparison), such as requiring access to the full history of client updates, which necessitates linking clients to their past updates and compromises the privacy design of federated learning. For fairness, we enforce full client participation, as required by PGA, even though our method naturally supports the standard partial participation setting. For FedQUIT, we always use one single epoch for unlearning ($E_{u}=1$) at the target client. We report the results for FedQUIT implemented as FedQUIT-Logits with $v=0$, as it performed best. In Sec. 6.1, we also discuss the results of FedQUIT-Softmax and other choices of $v$ value. See Appendix 8 for hyper-parameter tuning details.